OpenStudy (anonymous):

Can somebody please explain how to find the slope of the line determined by the following equation, without graphing it on paper? 3x=4y-2 I need a formula explained to be step by step, all the way to the end.

OpenStudy (anonymous):

Finding the slope is very simple. First, solve for y in terms of x. Then look at the coefficient on the x variable.. This is the slope.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

don't understand the 'solve for y in terms of x'...

OpenStudy (anonymous):

Basically you want to get y by itself on one side of the equal sign.

OpenStudy (anonymous):

I've tried just isolating y....did not get the answer I needed..

OpenStudy (anonymous):

\[ay + bx + c = 0 \implies y = \frac{-b}{a}x - \frac{c}{a}\]

OpenStudy (anonymous):

What did you get when you isolated the y?

OpenStudy (anonymous):

3/4x+1/2. The answer key to my college textbook says it's just 3/4.

OpenStudy (anonymous):

Right. Which is what I said.

OpenStudy (anonymous):

Once you isolate the y, you take the coefficient on the x variable. That is your slope.

OpenStudy (anonymous):

y = 3/4x+1/2 Therefore the coefficient on the x variable is 3/4 Therefore the slope is 3/4

OpenStudy (anonymous):

Got it now, thanks. :)

OpenStudy (amistre64):

the simplest way is just to determine 2 points from the equation; and find out how far the y moves when x moves 1 unit

OpenStudy (amistre64):

3x=4y-2 ; when x =0, y = 2/4 ; when x = 1, y = 5/4 the slope is the change in y; 5/4 - 2/4 = 3/4

OpenStudy (anonymous):

The amount the y moves when x moves one unit _is_ the coefficient on the x variable. ;p

OpenStudy (amistre64):

yes it is :)

OpenStudy (anonymous):

much easier to find then plugging in numbers ;p

OpenStudy (amistre64):

easier? I agree, but simplicity is not in moving and rearranging. Just another way to accomplish it ;)

OpenStudy (anonymous):

I've got to get back to my math now...thanks polpak. amistre....I use your method on other slope problems..but found it confusing on this one, glad that I now have a back-up formula. Thanks again. :)

OpenStudy (amistre64):

good luck ;)